Please explain, don't just give the answer!

28. Airway resistance is affected primarily by the A) amount of surfactant. B) thickness of the cartilage in the bronchial wall. C) amount of elastic tissue in the lungs. D) diameter of the bronchioles. E) partial pressure of each type of gas in inspired air
Answer #1

Airway resistance is affected primarily by the (D) diameter of the bronchioles

REASON:

Airway resistance =Pressure difference between atmospheric and alveolar pressure(cm H2O) / Airflow(L/sec)

Factors affecting airway resistance:

  1. Resistance to the airflow is inversely related to the total cross sectional area of the respiratory passage.When the diameter (cross section) is bigger,the resistance against the airflow is smaller and vice versa.Therefore,the resistance to the airflow is high in the conducting zone where as it is low in the respiratory zone
  2. According to Poiseuille-Hagen formula:

Equation(i) :Flow=

Pressure difference between two ends of the tube\times \left ( \pi /8 \right )\times \left ( 1/\eta \right )\times \left \left ( r^{4} /L\right )

\eta=Viscosity

L = Length of the tube

r=radius of the tube

Equation(ii): Since Flow=Pressure difference between two ends of the tube / Resistance(R)

Therefore from equation (i) and (ii),it follows

R=8\eta L/\pi r^{4}

R is directly proportional to \eta and L

R is inversely proportional to 1/r^{4}

Therefore,if the radius decreases by half,keeping the other factor constant the resistance increases by 16 times.